10. A new company is producing ant farms. They had a start-up fee of $200 and have supplies that cost $5 per ant farm. Write a linear equation that shows the cost for this new company creating a farms.
11. Using a table, find the cost to produce 8 farms, 10 farms and 12 farms. You will need to graph this below. Scale each axis.
12. Using your graph, find the cost to produce 20 and 22 of these ant farms.
13. What does the y-intercept in this specific problem tell you?
14. What does the slope indicate for this specific problem?

Answer by TimothyLamb(4379)

y = 5x + 200
---
cost for 8 farms:
y = 5*8 + 200
y = 240
---
cost for 10 farms:
y = 5*10 + 200
y = 250
---
cost for 12 farms:
y = 5*12 + 200
y = 260
---
cost for 20 farms:
y = 5*20 + 200
y = 300
---
cost for 22 farms:
y = 5*22 + 200
y = 310
---
y-intercept is the start-up cost ($200)
---
slope is the supply cost to produce one farm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php